Minutes of the code management group meeting 2015-05-06

From UKCA
Jump to: navigation, search

Code management group minutes


Present

Mohit Dalvi (MD), Coin Johnson (CJ), Nick Savage (NS), Fiona O'Connor (FOC), Nicolas Bellouin (NB), Luke Abraham (LA), Mark Richardson (MR), Graham Mann (GM), Janes Mulcahy (JM)

Apologies: Alex Archibald (AA)


Actions from this meeting

  1. JM to circulate GA7.0 email.
  2. JM/GM to talk offline regarding improvements
  3. JM to open ticket
  4. (ongoing) CJ to email LA re problems with SO2 diagnostics for emissions. Issues related to using ASAD diagnostics package.
  5. (ongoing) CJ to put branch on PUMA email CMG location of branch to check details
  6. (in progress) GM check 8.4 versus 7.3 GLOMAP-mode code and report any bug fixes the bugfix page
  7. (in progress) Alistair Sellar to add in status and UKESM1 ticket numbers to each item (re: UKESM1 strategy document)
  8. (ongoing) FOC/NB/GM - discuss (9) offline (re: GLOMAP-mode coupling to FJX) - Apostolos Voulgarakis could also be included in this. Could this also link to the heterogenous chemistry task?
  9. (in progress) ALL - feed in to Alistair Sellar list of priorities. Items need an owner if people need it raised.
  10. FOC to raise Rose stem with UM Systems Team, then LA may raise at UMPB or FOC with JWCRP
  11. LA to contact MONSooN team about Rose stem on MONSooN
  12. (not done) Incorporate Bug #35 into dry deposition work - Federico Centoni (FC) wants to do some sensitivities - hold off until FCs assessment - still waiting
  13. FOC to follow-up with FC
  14. (in progress) Bug #36 (ozone): FOC/Steven Hardiman to take on - see MOSRS ticket #102 (UKCA Bugfix for ozone column diagnostic calculation)
  15. NS to discuss wet_levels with Glenn Greed
  16. NS to possibly add-in run-time check for wet_levels for vn10.2, or remove from namelists and set to model_levels(?)
  17. (ongoing - standing item) CMG to monitor list of bugs and circulate between group for owners
  18. (ongoing - standing item) At CMG future meetings there should be included an appendix listing the open bugs. If a ticket is included it is not an open bug (as the owner of ticket is person doing lodging).
  19. (ongoing) ALL to migrate tickets over to the shared repository (MOSRS).
  20. (in progess) JM to link/send powerpoint results (re: GA7.0), updated to latest configuration and link/put files on UKCA wiki
  21. (in progress) MD to update suites to vn10.1
  22. NS to continue with lodging ticket #368 quick fix
  23. LA to contact Oliver Wild regarding NR solver
  24. (in progress) Bug #38: GM to take on lodging this on the trunk
  25. JM to contact Alistair Sellar and check on criticality of bug #38 for UKESM1.
  26. FOC to check documentation for emissions units
  27. (in progress) LA to check about NMVOC in new emissions routine
  28. LA to check doc for NVOC factor and open ticket to remove it from code
  29. (ongoing) ALL with access to keep monitoring ukca_bugs email list.
  30. (ongoing) ALL to think about if code changes will change KGO and let systems team know in good time.
  31. LA to wait until after Exec on 2015-05-08 to contact Jeremy Walton regarding vn7.3 on MONSooN Cray
  32. NS to contact Steve Mullerworth for information on Limited Area LFRic and UKCA should start to work with it
  33. FOC to raise with Ben Shipway (as well as MH from previous LFRic meeting) regarding when aerosols will be needed
  34. FOC to raise with David Walters about aerosol climatologies
  35. LA to open a doodle poll for 18th June to end of June.

Teleconference Numbers

  • UK Freefone: 0800 917 1956
  • Participant passcode: 16231039 then #

Agenda

  1. Minutes of the last meeing
  2. GA7.0 update
  3. Rose/Rose stem on MONSooN
  4. Bugfixes
  5. Targets for vn10.2
  6. LFRic
  7. A.O.B./D.N.M.
  8. Date of next meeting

Minutes

Minutes of the last meeing

Actions carried forward from previous meetings

  1. JM to email LA, cc GM (re: GA7.0/SO2) - done
    Action: JM to circulate email.
    Action: JM/GM to talk offline regarding improvements
    Action: JM to open ticket
  2. (ongoing) CJ to email LA re problems with SO2 diagnostics for emissions. Issues related to using ASAD diagnostics package. - ongoing
  3. (ongoing) CJ to put branch on PUMA email CMG location of branch to check details - ongoing
  4. (in progress) GM check 8.4 versus 7.3 GLOMAP-mode code and report any bug fixes to Mohit. - in progress(eye on GLOMAP-mode next version)
    Change of action: GM to now add to bugfixes page rather than emailing MD.
  5. (ongoing) FOC to add in status and ticket numbers to each item (re: UKESM1 strategy document) - in progress
    ownership transferred to Alistair Sellar who will add UKESM1 tickets. UM tickets done separately, and ticket numbers are recorded in UKESM1 tickets.
  6. (ongoing) FOC/NB/GM - discuss (9) offline (re: GLOMAP-mode coupling to FJX) - Apostolos Voulgarakis could also be included in this. Could this also link to the heterogenous chemistry task? - ongoing
  7. (ongoing) ALL - feed in to FOC list of priorities. Items need an owner if people need it raised. - in progress
    feedback now to Alistair Sellar
  8. LA to monitor NCAS-CMS plans for Rose stem, possibly ask for a subset - done
    Action: FOC to raise Rose stem with UM Systems Team, then LA may raise at UMPB or FOC with JWCRP
    Action: LA to contact MONSooN team about Rose stem on MONSooN
  9. (not done) Incorporate Bug #35 into dry deposition work - Federico Centoni (FC) wants to do some sensitivities - hold off until FCs assessment - still waiting
    Action: FOC to follow-up with FC
  10. (ongoing) Bug #36 (ozone): FOC/Steven Hardiman to take on - see MOSRS ticket #102 (UKCA Bugfix for ozone column diagnostic calculation) - in progress
    to be lodged at vn10.2.
  11. FOC/MD to open some tickets and add them to the bugfixes page, also there are some from Action/Task 4 from Previous Meetings which need adding. - done
  12. GM/MR to update bugfix page with their bug involving model_levels/wet_levels (will there even be a distinction in future, and/or could we just add a trap on this?) - done
  13. NS to check with Glenn Greed about wet_levels. - done
    Had extensive correspondance. Bottom line is wet_levels has never been implemented. Should be able to assume that wet_levels==model_levels. Don’t need to change current code, but future coding should use model_levels instead. It is Glenn Greed’s teams’ responsibility to do this UM-wide change. Same is also true of tracer_levels.
    Action: NS to discuss wet_levels with Glenn Greed
    Action: NS to possibly add-in run-time check for wet_levels for vn10.2, or remove from namelists and set to model_levels(?)
  14. (ongoing - standing item) CMG to monitor list of bugs and circulate between group for owners - ongoing
  15. (ongoing - standing item) At CMG future meetings there should be included an appendix listing the open bugs. If a ticket is included it is not an open bug (as the owner of ticket is person doing lodging).
  16. (ongoing) ALL to migrate tickets over to the shared repository (MOSRS). - ongoing
  17. JM to link/send powerpoint results (re: GA7.0) - in progress
    now a bit out of date from old presentation.
    Action: JM to update with latest config and put links/files to UKCA wiki.
  18. MD to email round MONSooN Rose suite information when working - done
  19. MD to update suites to vn10.1 - in progress
    MD is testing. MD thinks that MASS archiving may be fixed with branches.
  20. JM to check GA7.0 timeline - done
    has been nothing official. Still hasn’t been “chilled”. Likely that timeline will slip by at least a month. Will report when know. Setting up a vn10.0 job. TTL PEG would use vn10.0.
  21. MD to email AS and ask about netCDF emissions apps at vn10.1. - done
    Suite now uses netCDF emissions.
  22. Bug #31: NS to outline possible options and to open a ticket - done
    MOSRS ticket #368 opened. Gives 3 options. Convergence testing in NR doesn’t seem to be very clear.
    Action: NS to continue with lodging quick fix
    Action: LA to contact Oliver Wild
  23. Bug #38: GM to take on lodging this on the trunk - in progress
    Stretch for vn10.2 due to Rose metadata change, and more than just a one-line change. Criticality dependent on whether stratospheric aerosol configuration will be used for UKESM1.
    Action: JM to contact Alistair Sellar and check on criticality for UKESM1.
  24. Bug #41: FOC to follow this up and see if the netCDF code deals with this. - done
    Ben Johnson generated new ancillaries in kg(C), and has since updated these new ancillaries so that they are in kg(POM). NetCDF code seems to be correct. Bug #41 can be written off.
    Action: FOC to check documentation for emissions units
  25. Bug #41: LA to check about NMVOC in new emissions routine - in progress
    Action: LA to check doc for NVOC factor and open ticket to remove it from code
  26. ALL with access to keep monitoring ukca_bugs email list. - ongoing
  27. MR to send NS his vn8.6 branch. - done
  28. NS to contact MR about this work as it might impact him - done
  29. ALL to think about if code changes will change KGO and let systems team know in good time. - ongoing
  30. LA to ask Glenn Greed about the order in which tickets are committed. - done
    "no particular order"
  31. ALL to read the LFRic documents - done
  32. FOC to arrange doodle poll for discussion after Easter and arrange 1.5 hour meeting. - done
  33. FOC/LA to catch-up with Matt Hort - FOC to follow up. FOC/LA to update document accordingly. - done
  34. LA to check with Ros/Grenville exactly how hard it will be to get vn7.3 working on MONSooN-Cray - done
    UKESM team may take this on.
    Action: LA to wait until after Exec on 2015-05-08 to contact Jeremy Walton.
  35. LA to clarify which other UM versions are planned to be ported. - done
  36. LA to email all UKCA MONSooN users and highlight deadlines on data porting - done

GA7.0 update

Rose/Rose stem on MONSooN

Bugfixes

List of open bugs

Open bugs are classed as those with no associated ticket number and/or which are not classed as lodged.

  • #31 (FP exceptions in asad_sparse_vars or asad_spimjmp)
    Met Office ticket: #5616
    PUMA ticket: #555
    Comment: Temporary Work-around. May not work in all cases
    Previous Action: NS to outline possible options and to open a ticket.
  • #35 (Error in labelling: Surface resistance <--> stomatal resistance ?)
    Met Office ticket: #6417
    PUMA ticket: #653
    Comment: Awaiting input from Federico Centoni
    Previous Action: Holding off until assessment from Federico Centoni
  • #36 (Error in calculation of Column Ozone diagnostic)
    MOSRS ticket: #102
    Comment: To be lodged at vn10.2
    Previous Action: FOC to lodge
  • #38 (Missing logical for volcanic so2 emissions.)
    Met Office ticket: #6418
    PUMA ticket: #656
    Comment: To be Lodged
    Previous Action: GM to take on lodging this on the trunk
  • #41 (OC biomass burning emissions - incorrect conversion. Assumes emissions as kg POM, but (new) ancils are kg C)
    Met Office ticket: #6429
    Comment: To be lodged: Need to determine which ancils are POM vs C
    Note: netCDF emissions should explicitly deal with this, assuming that metadata is correct.
    Previous Action: FOC to follow this up and see if the netCDF code deals with this.
    Previous Action: LA to check about NMVOC in new emissions routine
  • #46 (Inconsistent use of wet_levels and model_levels in UKCA_AERO_CTL 1-D arrays)
    Met Office ticket: none
    Comment: Impact if wet_levels/= model_levels
    Previous Action: GM/MR to fix as part of other fixes
  • #47 (Dimension mismatch: q_chem, t_chem not considering extra level for ENDGame)
    Met Office ticket: #379 (MOSRS)
    Comment: To be lodged at UM10.2
  • #48 (Bugs in RADAER - dimensions, asymmetry param)
    Met Office ticket: #380 (MOSRS)
    Comment: To be lodged at UM10.2. Leads to div-by-zero error in rescale_phase_fnc
  • #49 (UKCA emissions: incorrect number, un-initialised slots in all_emissions())
    Met Office ticket: #381 (MOSRS)
    Comment: To be lodged at UM10.2
  • #50 (ukca_solve_coagnucl failures)
    Met Office ticket: #319 (MOSRS)
    Comment: Mainly a workaround. To be lodged at UM10.2

A lot will be combined into a package fix (MOSRS ticket #564). FOC will get in as many as possible.

Targets for vn10.2

Code review submission deadline is 12th June 2015 for vn10.2

LFRic

See also the LFRic discussion and draft UKCA-LFRic workplan circulated by LA.

Action: NS to contact Steve Mullerworth for information on Limited Area LFRic and UKCA should start to work with it

Action: FOC to raise with Ben Shipway (as well as MH from previous LFRic meeting) regarding when aerosols will be needed

Action: FOC to raise with David Walters about aerosol climatologies

Any other business/do not minute

Date of next meeing

Action: LA to open a doodle poll for 18th June to end of June.

See http://doodle.com/vyq2ti6m5v3b9zw8